Physics 115A xxxxx Fall 2013 Statistical Physics due 9:30 am Tuesday, October 29

PROBLEM SET 4

Reading: Chapter 5 in Reif. Midterm: Tuesday November 5 in class. The test will cover all material through Chapter 5 although we have not yet discussed all the applications of what is learned in Chapter 5. The exam will be closed book.

Problems:

Reif 5.1

Reif 5.2

Reif 5.4

Reif 5.7

Reif 5.10

Reif 5.12